OpenStudy (anonymous):

Use L'Hospitals Rule Limx->1 lnxtan(pi x/2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}lnx \tan Pix/2)\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}lnx *\tan(\pi x/2)\] \[\lim_{x \rightarrow 1} \frac{ lnx }{ \cot(\frac{ \pi x }{2 } )}\]

OpenStudy (anonymous):

I thought the d/dx of tan is sec^2x, no?

OpenStudy (anonymous):

oooh oh oh, i see

OpenStudy (anonymous):

I set it up wrong, its 1/tan. im not ready to take d/dx

OpenStudy (anonymous):

i put in this form (0/0)

OpenStudy (anonymous):

me jumping ahead of myself

OpenStudy (anonymous):

you can now use L'Hospitals Rule.

OpenStudy (anonymous):

right, ASAAD123 you are the best. thank you for all your help!!!

OpenStudy (anonymous):

welcome ^__-__^ ---